OpenStudy (anonymous):

help please,What is the radius of a circle with a circumference of 23.55 meters? Use 3.14 for pi and show your work.

OpenStudy (apoorvk):

Circumference = 2 x \(\pi\) x radius Can you use this to find the radius? Circumference and 'pi' values are already with you!

OpenStudy (anonymous):

So the only unknown you have is the radius....all you have to do it solve for it.

OpenStudy (anonymous):

Plug in all the values that you are given and solve for the radius

OpenStudy (anonymous):

23.55 m =2 x [\pi\] x r | /2 x pi 23.55 m /6.28 =r r = 3,75 m